# Guided Multi-Fidelity Bayesian Optimization for Data-driven Controller Tuning with Digital Twins

### v1

@@ III-C Adaptive Multi-Source Kernel Function
Building on the kernel formulation proposed for multi-fidelity GP models in [25, 26], we define the composite surrogate kernel as
@@
where $\mathcal{M}(r,l):=\left(1+\tfrac{\sqrt{5}r}{l}+\tfrac{5r^{2}}{3l^{2}}\right)\exp\!\left(-\tfrac{\sqrt{5}r}{l}\right)$ is the Matérn-5/2 kernel, $\sigma_{i}^{2}$ are output variances, and $p$ modulates correlation among sources with $s<1$. The hyperparameters $l_{i}$, $\sigma_{i}^{2}$, and $p$ are learned via MAP estimation by maximizing the marginal log-likelihood with log-priors on $l_{i}$ [22]. The mismatch metric $\hat{e}_{\text{IS2}}$ from (12) adjusts $\gamma_{0}$ to reduce IS2–IS1 correlation when IS2 accuracy degrades. Here, $c_{0}(k_{1},k_{2})$ captures correlation between IS1 and IS2 as a function of controller-parameter distance, while $\gamma_{1}$ captures correlations among IS2 and IS3 data, and multiplication by $c_{1}(k_{1},k_{2})$ further incorporates controller-parameter similarity. Overall, the kernel jointly accounts for distances in controller space and fidelity-dependent correlations across all information sources.
To adapt the IS2–IS1 correlation to the estimated IS2 accuracy, we use the clamping operator $\mathcal{P}_{[a,b]}(x):=\min\{b,\max\{a,x\}\}$, which projects $x\in\mathbb{R}$ into $[a,b]$. Let $s^{\prime}=\min(s_{1},s_{2})$. The cross–source lengthscale is defined as
$l_{\gamma_{0}}(\hat{e}_{\text{IS2}}):=\mathcal{P}_{[l_{\gamma_{0_{\text{min}}}},\,l_{\gamma_{0_{\text{max}}}}]}\!\left(\frac{s^{\prime}}{\hat{e}_{\text{IS2}}}\right),\vskip-2.39996pt$ (16)
where $\hat{e}_{\text{IS2}}$ is the normalized mismatch in (12). The clamp bounds $l_{\gamma_{0_{\text{min}}}},\,l_{\gamma_{0_{\text{max}}}}]:=[0.1,\,2]$ prevent under- or overfitting while keeping cross-source correlation responsive to $\hat{e}_{\text{IS2}}$. Large $\hat{e}_{\text{IS2}}$ (poor DT accuracy) pushes $l_{\gamma_{0}}$ toward $0.2$ (near-zero correlation), whereas small $\hat{e}_{\text{IS2}}$ (high DT accuracy) drives it toward $2$ (strong but not saturating correlation). We set $s^{\prime}=0.1$ as a heuristic to reflect the lower fidelity of the DT (IS2) relative to IS1. This choice also serves as a neutral pivot in (16): for $\hat{e}_{\text{IS2}}<s^{\prime}$, the lengthscale increases (stronger cross-source coupling), while for $\hat{e}_{\text{IS2}}>s^{\prime}$ it decreases (weaker coupling).
@@ III-D Adaptive Cost-aware Acquisition Function
We define the fidelity-aware sampling cost as
$\mathcal{H}(s|\hat{e}_{\text{IS2}}):=\begin{cases}1,&s=1,\\[4.0pt] \mathcal{P}_{[\mathcal{H}_{\text{min}},\,\mathcal{H}_{\text{max}}]}\bigl(\beta\cdot\hat{e}_{\text{IS2}}\bigr),&s\neq 1,\end{cases}\vskip-2.39996pt$ (17)
where $\beta\in\mathbb{R}$ is a constant scaling factor. Thus, IS1 evaluations always incur unit cost, while IS2 evaluations are scaled according to the estimated fidelity $\hat{e}_{\text{IS2}}$, with clamping to $\mathcal{H}_{\text{min}},\,\mathcal{H}_{\text{max}}]:=[0.1,1]$ ensuring numerical stability. We then define a cost-aware Expected Improvement acquisition function (caEI) as
$\text{caEI}_{n}(z|\hat{e}_{\text{IS2}}):=\frac{a_{\text{EI},n}(z)}{\mathcal{H}(s|\hat{e}_{\text{IS2}})},\vskip-2.39996pt$ (18)
where $a_{\text{EI},n}(z)$ is the expected improvement acquisition function with a closed-form formulation [27], computed independently for each information source. For fidelity $s$ at iteration $n$, we write
$a_{\text{EI},n}(z):=\sigma_{n}(z)\bigl(\upsilon_{n}(z)\Phi(\upsilon_{n}(z))+\varphi(\upsilon_{n}(z))\bigr),\vskip-2.39996pt$ (19)
with $\upsilon_{n}(z):=(\mu_{n}(z)-\mathsf{g}(s)^{+})\sigma_{n}(z)^{-1}$, where $\mu_{n}(z)$ and $\sigma_{n}(z)$ are the GP posterior mean and standard deviation, $\mathsf{g}(s)^{+}$ is the best observed objective in fidelity $s$, $\Phi(\cdot)$ is the standard normal cdf, and $\varphi(\cdot)$ is the pdf.
For mechatronic systems with first-principles models, we expect $\beta=4$ to be a reasonable starting point, so that high DT fidelity (small $\hat{e}_{\text{IS2}}$) yields values near 0.10.1—lowering the denominator of (18) and encouraging DT sampling—while low DT fidelity pushes the cost toward $1$, discouraging reliance on inaccurate simulations. For a given $\hat{e}_{\text{IS2}}$, increasing $\beta$ effectively decreases the sampling frequency from IS2; for instance, $\beta=10$ results in less IS2 queries. Clamping avoids degenerate cases: the lower bound prevents division by a minimal denominator, and the upper bound ensures DT queries are never more expensive than real-system ones. This formulation enables the optimizer to adaptively switch between IS1 and IS2 based on the current accuracy of the DT, thereby promoting data efficiency while maintaining robustness.
@@ III-E Guided Multi-Fidelity BO Algorithm
Our GMFBO framework leverages real data not to build a globally accurate digital twin (DT), but to make it objective-oriented: the DT is selectively corrected in regions of the parameter space that are relevant for reaching the optimization objective. This distinguishes our approach from classical active learning, which focuses on improving model accuracy, aiming to enhance predictive fidelity everywhere.
@@
1:Initialize $\mathcal{K}$, $N$, $N_{0_{s=1}}$, $N_{0_{s^{\prime}}}$, $n_{\text{c}}$, $\mathcal{D}_{\text{c}}$, $\mathrm{GP}_{\text{c}}$, $\mathcal{D}\leftarrow\mathcal{D}_{\text{init}}$, $\hat{e}_{\text{IS2}}$, $s^{\prime}$, $\alpha$, $\beta$
@@
3: Train surrogate GP model given $\mathcal{D}$
@@
6:  Evaluate real system: $(t,\mathrm{y}(k,t,s),\mathsf{g}(k,s))\leftarrow\text{IS1}(k)$
7:  Update $\mathcal{D}\leftarrow\mathcal{D}\cup\{[k,1],\mathsf{g}(k,s)\}$
8:  Evaluate DT: $\mathrm{y}(k,t,s^{\prime})\leftarrow\text{IS2}(k)$
9:  Update $\mathcal{D}_{\text{c}}\leftarrow\{([t,\mathrm{y}(k,t,s^{\prime})],\mathrm{y}(k,t,1))\,|\,t\in[0,T]\}$
10:  Update correction model $\mathrm{GP}_{\text{c}}$
11:  while $n_{\text{c}}$ new data from IS3 is not added to $\mathcal{D}$ do
12:   Draw $k^{\prime}\in\mathcal{K}$
13:   Evaluate corrected DT: $\mathsf{g}_{\text{c}}(k^{\prime},s^{\prime})\leftarrow\text{IS3}(k)$
14:   if (10) is satisfied then
15:    Update $\mathcal{D}\leftarrow\mathcal{D}\cup\{[k^{\prime},s^{\prime}],\mathsf{g}_{\text{c}}(k^{\prime},s^{\prime})\}$
16:   end if
17:  end while
18:  Update simulation fidelity estimation: $\hat{e}_{\text{IS2}}$
19:  Update $l_{\gamma_{0}}$ and $\mathcal{H}$ based on new $\hat{e}_{\text{IS2}}$
20:else if $s=s^{\prime}$ then
21:  Evaluate DT: $\mathsf{g}(k,s^{\prime})\leftarrow\text{IS2}(k)$
22:  Update $\mathcal{D}\leftarrow\mathcal{D}\cup\{[k,s^{\prime}],\mathsf{g}(k,s^{\prime})\}$
23:end if
24:end for
25:return $k^{*}\leftarrow\arg\min\limits_{k}\mathsf{g}(k,s=1)\,,\>\forall(k,1)\in\mathcal{D}$
@@ IV System Modelling
We employ the high-efficiency joint (HEJ) 90 drive system from maxon to iteratively tune the parameters of its joint position-velocity-torque (JPVT) controller. The system comprises a high-torque permanent magnet synchronous motor, a planetary gearbox, dual encoders, and power electronics, with an EPOS4 controller that features a hierarchical structure consisting of a low-level current loop and higher-level position, velocity, and torque loops. The JPVT controller uses a proportional–derivative position strategy with tunable parameters $k=(K_{p},K_{d})$, constrained by
@@
Fig. 3(a) shows the information-source block diagram. Our objective is a weighted sum of four performance metrics from the position step response. A unit step reference $\mathrm{y}^{*}(t)=0$ for $t<t_{0}=0.1$ [s] and $\mathrm{y}^{*}(t)=1$ [rad] ($652$ encoder counts) for $0\leq t-t_{0}\leq T=1$ [s] is applied to the motor output, yielding $\alpha=1$ from (10). We define $\mathrm{h}(k)=[O_{\text{s}},T_{\text{tr}},T_{\text{r}},T_{\text{s}}]$, where $O_{\text{s}}$, $T_{\text{tr}}$, $T_{\text{r}}$, and $T_{\text{s}}$ denote overshoot, transient, rise, and settling time, respectively. The objective function $\mathsf{g}$ for augmented input $z$ is
$\mathsf{g}(z)=w_{1}\cdot O_{\text{s}}+w_{2}\cdot T_{\text{tr}}+w_{3}\cdot T_{\text{r}}+w_{4}\cdot T_{\text{s}},\vskip-1.20007pt$ (21)
with weights $w_{i},\,i=1,\ldots,4$. Since $\hat{e}_{\text{IS2}}$ in (12) is based on these metrics, the mismatch mainly reflects transient step-response errors, while effects outside this window or uncaptured by the metrics (e.g., backlash) may remain unobserved. To construct $\mathcal{D}_{\text{init}}$, we randomly sample $(K_{p},K_{d})$ pairs from $\mathcal{K}$ and record the resulting $\mathrm{h}(k)$. The $\mathrm{GP}_{\text{c}}$ model is retrained whenever $\mathcal{D}_{\text{c}}$ is updated. Fig. 3(b) illustrates IS3, where the learned model corrects DT outputs.
@@
We employ a digital twin of the drive system (IS2), based on a nonlinear inverse flux linkage motor model [30] refined to account for friction, ripple harmonics, and a simple gearbox backlash model [31]. Table I summarizes the chosen values for the design parameters of our method.
@@ V Numerical Analysis
Unless stated otherwise, we set $N_{0_{s=1}}=2$, $N_{0_{s^{\prime}}}=10$, and $N=20$ BO iterations. The observation noise variance $\sigma_{\eta}^{2}$ is shared across information sources and estimated via GP marginal log-likelihood [32]. We use $n_{\text{c}}=4$ corrected IS3 estimates and optimize the cost-aware EI in (18) with L-BFGS [33] to select the next candidate $(k_{n+1},s_{n+1})$. Objective weights and normalization are computed from $10$ Latin Hypercube samples (LHS) [34], yielding $w_{1},w_{2},w_{3},w_{4}]=[0.02,0.20,0.70,0.20]$, with means and standard deviations used for normalization of objectivess . Each tuning run is repeated $N_{\text{exper}}=50$ times with $\mathcal{D}_{\text{init}}$ from random LHS samples. Let $n^{*}$ be the average IS1 iterations to reach the optimum.
Fig. 4(a) shows the ground-truth objective over $\mathcal{K}$, where optimal regions occur at larger gains. Fig. 4(b) shows IS2 estimates with $20\%$ relative error in $60\%$ of $\mathcal{K}$, but $50\%$ error in over $10\%$ of the domain. Errors are largest away from the optimum, enabling the correction GP to refine DT accuracy in the optimal region and more effectively guide the optimizer.
Figure 4: (a) True objective function given IS1 and (b) relative absolute error (in %) of the estimated objective by IS2
V-A Initial Dataset Ablation Study
Table II reports the required IS1 iterations ($n^{*}$) for different initial datasets. GMFBO performs best with a balanced amount of IS2 data: excess data degrades the surrogate due to IS2 inaccuracies, while insufficient data reduces efficiency. The benefit is most pronounced when IS1 samples are scarce.

### v2

@@ III-C Adaptive Multi-Source Kernel Function
We describe how different data sources are coupled within the multi-fidelity GP surrogate model for the BO objective. Building on the kernel formulation proposed for multi-fidelity GP models in [25, 26], we define the composite surrogate kernel for the BO optimization objective model as
@@
where $\mathcal{M}(r,l):=\left(1+\tfrac{\sqrt{5}r}{l}+\tfrac{5r^{2}}{3l^{2}}\right)\exp\!\left(-\tfrac{\sqrt{5}r}{l}\right)$ is the Matérn-5/2 kernel, $\sigma_{i}^{2}$ are output variances, and $p$ modulates correlation among sources with $s<1$. The hyperparameters $l_{i}$, $\sigma_{i}^{2}$, and $p$ are learned via MAP estimation by maximizing the marginal log-likelihood with log-priors on $l_{i}$ [23]. The mismatch metric $\hat{e}_{\text{IS2}}$ from (9) adjusts $\gamma_{0}$ to reduce cross-fidelity correlation when IS2 accuracy degrades. Here, $c_{0}(k_{1},k_{2})$ captures correlation between IS1 (high fidelity) and IS2 or IS3 (low fidelity) as a function of controller-parameter distance, while $\gamma_{1}$ captures correlations between IS2 and IS3 data (both low fidelity sources), and multiplication by $c_{1}(k_{1},k_{2})$ further incorporates controller-parameter similarity. Overall, the kernel jointly accounts for distances in controller space and fidelity-dependent correlations across all information sources.
To adapt the IS2–IS1 correlation to the estimated IS2 accuracy, we use the clamping operator $\mathcal{P}_{[a,b]}(x):=\min\{b,\max\{a,x\}\}$, which projects $x\in\mathbb{R}$ into $[a,b]$. The cross–source lengthscale is defined as
$l_{\gamma_{0}}(\hat{e}_{\text{IS2}}):=\mathcal{P}_{[l_{\gamma_{0_{\text{min}}}},\,l_{\gamma_{0_{\text{max}}}}]}\!\left(\frac{1}{\hat{e}_{\text{IS2}}}\right),\vskip-2.39996pt$ (13)
where $\hat{e}_{\text{IS2}}$ is the mismatch in (9). The clamp bounds $l_{\gamma_{0_{\text{min}}}},\,l_{\gamma_{0_{\text{max}}}}]:=[0.1,\,1]$ prevent under- or overfitting while keeping cross-source correlation responsive to $\hat{e}_{\text{IS2}}$. Large $\hat{e}_{\text{IS2}}$ (poor DT accuracy) pushes $l_{\gamma_{0}}$ toward $0.1$ (near-zero correlation), whereas small $\hat{e}_{\text{IS2}}$ (high DT accuracy) drives it toward $1$, enabling moderate correlation while preventing overreliance on simulation data.
@@ III-D Adaptive Cost-aware Acquisition Function
The information source for the next measurement is selected by extending the expected improvement (EI) acquisition to incorporate source-dependent sampling cost. We define the fidelity-aware sampling cost as
$\mathcal{H}(s|\hat{e}_{\text{IS2}}):=\begin{cases}1,&s=1,\\[4.0pt] \mathcal{P}_{[\mathcal{H}_{\text{min}},\,\mathcal{H}_{\text{max}}]}\bigl(\hat{e}_{\text{IS2}}\bigr),&s=0.\end{cases}\vskip-3.60004pt$ (14)
IS1 evaluations always have unit cost, while the cost for IS2 evaluations increases with the mismatch $\hat{e}_{\text{IS2}}$, with clamping to $\mathcal{H}_{\text{min}},\,\mathcal{H}_{\text{max}}]:=[0.1,1]$ ensuring numerical stability. We then define a cost-aware Expected Improvement acquisition function (caEI) as
$\text{caEI}(z|\hat{e}_{\text{IS2}}):=\frac{a_{\text{EI}}(z)}{\mathcal{H}(s|\hat{e}_{\text{IS2}})},\vskip-2.39996pt$ (15)
where $a_{\text{EI}}(z)$ is the EI acquisition function with a closed-form formulation [27]. It is computed independently for each information source such that
$a_{\text{EI}}(z):=\sigma(z)\bigl(\upsilon(z)\Phi(\upsilon(z))+\varphi(\upsilon(z))\bigr),\vskip-2.39996pt$ (16)
with $\upsilon(z):=(\mu(z)-\mathsf{g}(s)^{+})\sigma(z)^{-1}$, where $\mu(z)$ and $\sigma(z)$ are the surrogate GP posterior mean and standard deviation, $\mathsf{g}(s)^{+}$ is the best observed objective in fidelity $s$, $\Phi(\cdot)$ is the standard normal cdf, and $\varphi(\cdot)$ is the pdf. When DT fidelity is high (small mismatch $\hat{e}_{\text{IS2}}$), the denominator in (15) decreases, encouraging DT sampling due to increasing caEI. When $\hat{e}_{\text{IS2}}$ is high, the cost approaches 1, discouraging reliance on inaccurate simulations. IS3 points are added after IS1 is selected, not through the acquisition function. Clamping avoids degenerate cases: the lower bound prevents division by a minimal denominator, and the upper bound ensures DT queries are never more expensive than real-system ones. This formulation enables the optimizer to adaptively switch between IS1 and IS2 based on the current accuracy of the DT.
@@ III-E Guided Multi-Fidelity BO Algorithm
Our GMFBO framework leverages real data to refine the DT and provide more accurate performance estimates to guide the optimizer. The DT is selectively corrected in regions of the parameter space that are relevant for reaching the optimization objective. This distinguishes our approach from classical active learning, which focuses on improving model accuracy, aiming to enhance predictive fidelity everywhere.
@@
1:Initialize $\mathcal{K}$, $N$, $n_{\text{c}}$, $\mathfrak{D}_{\text{c}}$, $\mathrm{GP}_{\text{c}}$, $\mathfrak{D}\leftarrow\mathfrak{D}_{\text{init}}$, $\hat{e}_{\text{IS2}}$, $\alpha$
@@
3: Train surrogate GP model given $\mathfrak{D}$
@@
6:  Evaluate real system: $(t,\mathrm{y}(t,k,1),\mathsf{g}(k,1))\leftarrow\text{IS1}(k)$
7:  Update $\mathfrak{D}\leftarrow\mathfrak{D}\cup\{[k,1],\mathsf{g}(k,1)\}$
8:  Evaluate DT: $\mathrm{y}(t,k,0)\leftarrow\text{IS2}(k)$
9:  Update $\mathfrak{D}_{\text{c}}\leftarrow\{([t,\mathrm{y}(t,k,0)],\mathrm{y}(t,k,1))\,|\,t\in[0,T]\}$
10:  Update correction model $\mathrm{GP}_{\text{c}}\;,\;i\leftarrow 0$
11:  while $i<n_{\text{c}}$ do
12:   Sample $k^{\prime}\sim\mathcal{N}(k,\,0.05\,\mathbb{K})\in\mathcal{K}$
13:   Evaluate corrected DT: $\mathsf{g}_{\text{c}}(k^{\prime},0)\leftarrow\text{IS3}(k^{\prime})$
14:   Compute $\bar{\sigma}_{c}$ from (7)
15:   if $\bar{\sigma}_{\text{c}}\ll\alpha$ then
16:    Update $\mathfrak{D}\leftarrow\mathfrak{D}\cup\{[k^{\prime},0],\mathsf{g}_{\text{c}}(k^{\prime},0)\}\;,\;i\leftarrow i+1$
17:   end if
18:  end while
19:  Recompute the mismatch $\hat{e}_{\text{IS2}}$ from (9)
20:  Update $l_{\gamma_{0}}$ (12) and $\mathcal{H}$ (14) based on new $\hat{e}_{\text{IS2}}$
21:else if $s=0$ then
22:  Compute DT performance: $\mathsf{g}(k,0)\leftarrow\text{IS2}(k)$
23:  Update $\mathfrak{D}\leftarrow\mathfrak{D}\cup\{[k,0],\mathsf{g}(k,0)\}$
24:end if
25:end for
26:return $k^{*}\leftarrow\arg\min\limits_{k}\mathsf{g}(k,s=1)\,,\>\forall(k,1)\in\mathfrak{D}$
@@ IV System Modelling
We use the high-efficiency joint (HEJ) 90 drive system from maxon, and iteratively tune the parameters of its joint position-velocity-torque (JPVT) controller. The system comprises a high-torque permanent magnet synchronous motor, a planetary gearbox, dual encoders, and power electronics, with an EPOS4 controller that features a hierarchical structure consisting of a low-level current loop and higher-level position, velocity, and torque loops. The JPVT controller uses a proportional–derivative position strategy with tunable parameters $k=(K_{p},K_{d})$, constrained by
@@
Figure 3(a) shows the information-source block diagram. Our objective is a weighted sum of four performance metrics from the position step response. A unit step reference $\mathrm{y}^{*}(t)=0$ for $t<t_{0}=0.1$ [s] and $\mathrm{y}^{*}(t)=1$ [rad] ($652$ encoder counts) for $0\leq t-t_{0}\leq T=1$ [s] is applied to the motor output, yielding $\alpha=1$ from (8). We define $\mathrm{h}(k)=[O_{\text{s}},T_{\text{tr}},T_{\text{r}},T_{\text{s}}]$, where $O_{\text{s}}$, $T_{\text{tr}}$, $T_{\text{r}}$, and $T_{\text{s}}$ denote overshoot, transient, rise, and settling time, respectively. The objective function $\mathsf{g}$ for augmented input $z$ is
$\mathsf{g}(z)=w_{1}\cdot O_{\text{s}}+w_{2}\cdot T_{\text{tr}}+w_{3}\cdot T_{\text{r}}+w_{4}\cdot T_{\text{s}},\vskip-2.39996pt$ (18)
with weights $w_{i},\,i=1,\ldots,4$. Since $\hat{e}_{\text{IS2}}$ in (9) is based on these metrics, the mismatch mainly reflects transient step-response errors, while effects outside this window or uncaptured by the metrics (e.g., backlash) may remain unobserved. To construct $\mathfrak{D}_{\text{init}}$, we randomly sample $(K_{p},K_{d})$ pairs from $\mathcal{K}$ and record the resulting $\mathrm{h}(k)$. The $\mathrm{GP}_{\text{c}}$ model is retrained whenever $\mathfrak{D}_{\text{c}}$ is updated. Figure 3(b) illustrates IS3, where the learned model corrects DT outputs.
@@
We use a digital twin of the drive system (IS2), based on a nonlinear inverse flux linkage motor model [30] refined to account for friction, ripple harmonics, and a simple gearbox backlash model [31]. Table I summarizes the chosen values for the design parameters of our method.
@@ V Numerical Analysis
Unless stated otherwise, we set $N_{0_{s=1}}=2$, $N_{0_{s=0}}=10$, and $N=20$ BO iterations. A shared observation noise variance $\sigma_{\eta}^{2}$ across information sources is estimated via GP marginal log-likelihood [32]. We use $n_{\text{c}}=4$ IS3 corrections and optimize the cost-aware EI in (15) using L-BFGS [33] to select $(k_{n+1},s_{n+1})$. Objective weights and normalization are computed from $10$ Latin Hypercube samples (LHS) [34], yielding $w_{1},w_{2},w_{3},w_{4}]=[0.02,0.20,0.70,0.20]$, with means and standard deviations used for normalization of objectives. We perform a Monte Carlo study with $N_{\text{exper}}=50$ experiments, each initialized with $\mathfrak{D}_{\text{init}}$ drawn from random LHS.
Figure 4 shows the ground-truth objective over $\mathcal{K}$ (top), and the relative prediction error of IS2 (bottom). Errors exceed $100\%$ in some regions, causing IS2 data to mislead the optimizer by distorting the GP surrogate. This motivates correcting IS2 predictions via IS3 and adapting the surrogate and acquisition function to account for variable simulation fidelity.
Figure 4: The top panel shows the true objective function given IS1; The bottom panel shows the relative absolute error (in %) of the estimated objective by IS2
V-A Comparison with Baseline Methods
We compare GMFBO against baseline BO with EI and three MFBO variants: baseline MFBO with trace-aware Knowledge Gradient (taKG) [35], baseline MFBO with our caEI, and a modified MFBO with caEI using the fidelity-aware kernel in (10) without including IS3. The taKG acquisition weighs expected information gain against sampling cost, which relies strongly on the surrogate GP. GMFBO explicitly corrects DT bias and guides optimization toward the objective. In contrast, taKG resembles an active learning strategy that improves surrogate accuracy in informative regions without adapting to biased or shifting fidelities.
We first conduct an ablation study on the size of initial data per information source in $\mathfrak{D}_{\text{init}}$. Let $n^{*}$ denote the average IS1 iterations to reach the optimum. Table II reports $n^{*}$ with different initial datasets. GMFBO performs best with a balanced amount of IS2 data: excess data degrades the surrogate due to IS2 inaccuracies, while insufficient data reduces efficiency. The benefit is most pronounced when IS1 samples are scarce.